What is the value of K for this aqueous reaction at 298 K? Delta G = 19.54 kj/mol
Answer #1

Given data,

Temperature = 298K

ΔG = 19.54 KJ/mol

= 19540 J/mol

Let us consider a reaction,

ΔG = -R * T * ln Kc

19540 = - 8.314 * 298.0 * ln(Kc)

ln Kc = - 19540 / 2477.572

ln Kc = -7.88

Kc = e-7.88

Kc = 3.78 * 10-4